# Operator Spreading and the Absence of Many-Body Localization

### v2

@@ Appendix A Localized Models
We now denote by $\tilde{S}(k,j)$ our ‘adjusted’ Stirling number which inserts the requirement that contributions from sites of increasing distance from our initial site are exponentially decaying. Let $k_{n}$ denote the number of bonds at a distance $n$ from our initial site, where a distance of zero corresponds to a bond involving our initial site. We want exponential decay of the ratio of the contributions from bonds at $k_{n}$ to bonds at $k_{0}$. To insert this condition, we multiply by a factor of $e^{-\xi nk_{n}}$ for each $k_{n}$ and for any constant $\xi>0$. We note that this is normalized so that at $k_{0}$ this factor is 1, meaning the initial site is not reduced. This separate treatment of the $k_{n}$’s, which represent the number of times each bond appears in the connected cluster, is why the form (A.1) of $S(k,j)$ is desirable. Applying the decaying factors, we get
@@
$=$ $\sum_{k_{0}+\ldots+k_{j-1}=k}\frac{1}{j!}\prod_{i}\frac{1}{a_{i}!}{{k}\choose{k_{0},\ldots,k_{j-1}}}\prod_{n=0}^{j-1}e^{-\xi nk_{n}}$
$\leq$ $\sum_{k_{0}+\ldots+k_{j-1}=k}{{k}\choose{k_{0},\ldots,k_{j-1}}}\prod_{n=0}^{j-1}e^{-\xi nk_{n}}$
$=$ $\left(\frac{e^{\xi}(1-e^{-\xi j})}{e^{\xi}-1}\right)^{k}.$
We also loosened the condition on the $k_{i}$’s in the second line so that they are now greater than or equal to 00 instead of 1 so that we may apply the multinomial theorem in the third line. Since this may only increase the sum due to allowing more solutions to $k_{0}+\ldots+k_{j-1}=k$, the upper bound is still valid. Then the adjusted number of connected clusters, with the inserted requirement of the exponentially decreasing contribution with cluster length, is bounded from above by
$\sum_{\{I_{1},\cdots,I_{k}\}}\!\!\!\!\!\!\!{}^{\prime}$ $=$ $\sum_{j=1}^{k}2^{j}\tilde{S}(k,j)$ (A.5)
$\leq$ $2(1-e^{-\xi k})^{k}\left(\frac{e^{\xi}}{e^{\xi}-1}\right)^{k}(2^{k}-1)$
$\approx$ $2\left(\frac{2e^{\xi}}{e^{\xi}-1}\right)^{k}$
where in the last line we took the large $k$ limit. Defining the constant $\alpha=\frac{2e^{\xi}}{e^{\xi}-1}$, we therefore obtain the bound
@@
Appendix B Graphs
@@ B.1 No disorder
$p(m,k)$ $=$ $\sum_{i_{1}+\ldots+i_{m}=\frac{k}{2}}\;\prod_{j=1}^{m}C(i_{j}-1)$ (B.2)
@@
$s^{(1)}_{1}(k)=4^{k}\sum_{m=1}^{\frac{k}{2}}2^{m}\frac{m}{k-m}{{k-m}\choose{\frac{k}{2}}}=4^{k}{{k}\choose{\frac{k}{2}}}.$ (B.3)
@@
$s^{(1)}_{2}(k)=\frac{1}{4}s_{1}(k+1)=2^{2k+1}{{k}\choose{\frac{k-1}{2}}}\,.$ (B.4)
@@
$s^{(1)}_{l}(k)=\frac{1}{4}s_{l-1}(k+1)-s_{l-2}(k).$ (B.6)
Simple inductive arguments show that the solution to Eq. (B.6) with base cases  (B.4) and  (B.1) is indeed our previous result  (III.1.1).
B.2 Disordered case
@@
$||[H,\sigma^{z}_{0}]||_{1}\leq 8^{k}(D+1)^{k-1}.$ (B.7)
This is the result Eq. (III.14). The bound is exponential as expected, with an increasing disorder strength corresponding to an increased growth rate. We also see that setting $D=0$ gives us $8^{k}$, the exact result in the case without disorder.
@@
where ${}_{2}F_{1}$ is the hypergeometric function. Note that this result is derived for even $k$. For $k$ odd the support will be zero; one can infer this from Fig. 4 since there is no possible walk on the graph beginning and ending at the top node with an odd number of steps. The upper bound in Eq. (B.2) only arises from the fact that $D$ is an upper bound on the local magnetic fields; the enumerative calculations are exact. Similar numerical calculations show that the support on $l>1$ sites is bounded from above by a hypergeometric function as well but with a different leading term
@@
This coincides with Eq. (III.17). As mentioned in the main text, there is a maximum in this ratio which shifts to larger $l$ with increasing $k$, which is not expected for a localized model. This is likely explained by the looseness of the bound Eq. (B.2); while it is correct for the leading term in $D$, the actual prefactors for terms of decreasing order in $D$ become increasingly smaller than those in the given bound. We expect that a formula replacing Eq. (B.2) in which the enumerative calculations are exact and the upper bound only comes from $D$ would not possess this unexpected behaviour, although we have been unable to derive such a formula. We nevertheless see the expected result that with $k$ held constant we have an exponential decay of the ratio of support for increasing $l$, and that increasing $D$ corresponds to an increasing rate of decay.

### v3

@@ Appendix A Localized Models
We now denote by $\tilde{S}(k,j)$ our ‘adjusted’ Stirling number which inserts the requirement that contributions from sites of increasing distance from our initial site are exponentially decaying. Let $k_{n}$ denote the number of bonds at a distance $n$ from our initial site, where a distance of zero corresponds to a bond involving our initial site. We want exponential decay of the ratio of the contributions from bonds at $k_{n}$ to bonds at $k_{0}$. To insert this condition, we multiply by a factor of $e^{-nk_{n}/\xi}$ for each $k_{n}$ and for any constant $\xi>0$ which plays the role of a localization length. We note that this is normalized so that at $k_{0}$ this factor is 1, meaning the initial site is not reduced. This separate treatment of the $k_{n}$’s, which represent the number of times each bond appears in the connected cluster, is why the form (A.1) of $S(k,j)$ is desirable. Applying the decaying factors, we get
@@
$=$ $\sum_{k_{0}+\ldots+k_{j-1}=k}\frac{1}{j!}\prod_{i}\frac{1}{a_{i}!}{{k}\choose{k_{0},\ldots,k_{j-1}}}\prod_{n=0}^{j-1}e^{-nk_{n}/\xi}$
$\leq$ $\sum_{k_{0}+\ldots+k_{j-1}=k}{{k}\choose{k_{0},\ldots,k_{j-1}}}\prod_{n=0}^{j-1}\frac{e^{-nk_{n}/\xi}}{n+1}$
$=$ $\text{e}^{k/\xi}\left(\sum_{n=1}^{j}\frac{\text{e}^{-n/\xi}}{n}\right)^{k}.$
We also loosened the condition on the $k_{i}$’s in the second line so that they are now greater than or equal to 00 instead of 1 so that we may apply the multinomial theorem in the third line. Since this may only increase the sum due to allowing more solutions to $k_{0}+\ldots+k_{j-1}=k$, the upper bound is still valid. For $k\gg j\gg 1$ we can extend the sum up to infinity and obtain
$\tilde{S}(k,j)\leq\left(-\text{e}^{1/\xi}\ln(1-\text{e}^{-1/\xi})\right)^{k}\,.$ (A.5)
Then the adjusted number of connected clusters, with the inserted requirement of the exponentially decreasing contribution with cluster length, is bounded from above by
$\sum_{\{I_{1},\cdots,I_{k}\}}\!\!\!\!\!\!\!{\vphantom{\sum}}^{\prime}$ $=$ $\sum_{j=1}^{k}2^{j}\tilde{S}(k,j)$ (A.6)
$\leq$ $2(2^{k}-1)\left(-\text{e}^{1/\xi}\ln(1-\text{e}^{-1/\xi})\right)^{k}$
$\approx$ $2\left(-2\text{e}^{1/\xi}\ln(1-\text{e}^{-1/\xi})\right)^{k}$
where in the last line we took the large $k$ limit. Defining the constant $\alpha=-2\text{e}^{1/\xi}\ln(1-\text{e}^{-1/\xi})$, we therefore obtain the bound
@@
Appendix B Lieb-Robinson bounds
Another physically relevant property which can be obtained from our results for the $k$-fold commutator are Lieb-Robinson bounds for the spreading of information in Euclidean time.
B.1 General bound
We start by summarizing the result for the generic, quantum chaotic case which was first derived in Ref. [30]. We consider the commutator between two local operators a distance $\ell$ apart
$||[A_{0}(\tau),B_{\ell}(0)]||$ $=$ $||[\sum_{k}[H,A_{0}]^{(k)}\frac{\tau^{k}}{k!},B_{\ell}]||$
$\leq$ $2||A||\,||B||\sum_{k}\frac{(2CJ\tau)^{k}}{k!}\sum_{j=\ell}^{k}2^{j}S(k,j)$
where we have used the fact that the lattice animal has to consist of at least $j=\ell$ bonds for the commutator to be non-zero. We can now use the representation of the Stirling numbers of the second kind
$S(k,j)=\sum_{s=1}^{j}\frac{(-1)^{j-s}s^{k-1}}{(j-s)!(s-1)!}$ (B.2)
to rewrite the bound of the commutator as
$2||A||\,||B||\sum_{j=\ell}^{\infty}2^{j}\sum_{s=1}^{j}\frac{(-1)^{j-s}}{s(j-s)!(s-1)!}\underbrace{\sum_{k=1}^{\infty}\frac{(2CJ\tau s)^{k}}{k!}}_{\text{e}^{2CJ\tau s}-1}$
$=2||A||\,||B||\sum_{j=\ell}^{\infty}\frac{(2q)^{j}}{j!},\quad q=\text{e}^{2CJ\tau}-1$ (B.3)
[…]
For $\ell\gg q$ we obtain the scaling
$||[A_{0}(\tau),B_{\ell}(0)]||$ $\sim$ $2||A||\,||B||\frac{(2q)^{l}}{l!}$
$\sim$ $2||A||\,||B||(2q)^{\ell}\left(\frac{e}{\ell}\right)^{\ell}\,.$
This means that the commutator is $\mathcal{O}(1)$ if $q\sim\ell$ which is equivalent to $\exp(2CJ\tau)\sim\ell$. I.e., in a generic, chaotic one-dimensional quantum system with short-range interactions, information is spreading exponentially fast in Euclidean time. From the derivation we see that this exponentially fast spreading is directly related to the almost factorial growth of the commutator $[H,A]^{(k)}$.
B.2 Non-interacting bound
Let us first consider the case without disorder. In this case we have an explicit result for the contribution of lattice animals with $j$ bonds, see Eq. (III.1.1). Generalizing this to the case of an operator $A$, we find
$||[A_{0}(\tau),B_{\ell}(0)]||\leq 2||A||\,||B||\sum_{k=\ell-1}^{\infty}\frac{(4\tau)^{k}}{k!}\sum_{j=\ell}^{k+1}\binom{k}{\frac{k+1-j}{2}}$
$=2||A||\,||B||\sum_{k=\ell-1}^{\infty}\frac{(4\tau)^{k}}{k!}\sum_{j=0}^{(k+1-\ell)/2)}\binom{k}{j}$ (B.5)
for the case where $k$ is odd, and $\ell$ is even. This becomes non-zero if $k\gtrsim\ell$ and is of order $\mathcal{O}(1)$ if $\tau\sim\ell$. I.e., for a non-interacting system the spreading is ballistic in Euclidean time. This is directly related to the exponential instead of almost factorial growth of $[H,A]^{(k)}$ which is due to the fact that most lattice animals are zero.
Importantly, this scaling remains intact in the non-interacting disordered case. I.e., the spreading in Euclidean time is still ballistic but the time scale is renormalized. This can be seen by considering
$||[A_{0}(\tau),B_{\ell}(0)]||\leq 2||A||\,||B||\!\!\!\sum_{k=\ell-1}^{\infty}\frac{(\tau)^{k}}{k!}\sum_{j=\ell}^{k+1}\underbrace{||[H,A_{0}]^{(k)}||}_{2^{2j-1}\eta^{k-j+1}}$ (B.6)
where we have used that $s^{(1)}_{j}(k)=||[H,A_{0}]_{j}^{(k)}||\sim\theta(k-(j-1))2^{2j-1}\eta^{k-j+1}$. I.e., independent of the disorder we always have that $s^{(1)}_{j}(j-1)=2^{2j-1}$, see Eq. (III.13). Only the further increase of $s^{(1)}_{j}(k)$ for $k\geq j-1$ is affected by disorder, $\eta=\eta(D)$. However, this does not affect the scaling when the commutator becomes of order $\mathcal{O}(1)$ which is still given by $\tau\sim\ell$.
B.3 Many-body localization
In the interacting, disordered case all the paths which exist in the non-disordered case remain present, see Fig. 13. This suggests that information will still spread exponentially fast, $\text{e}^{\tau}\sim\ell$, in Euclidean time albeit with a renormalized time scale. I.e., based on the view of the commutator $[H,A_{0}]^{(k)}$ as a graph we expect that while the spreading of information does slow down with increasing disorder, this is merely a rescaling of time but the functional form remains unaffected.
The only way how the spreading of information can change qualitatively is if disorder introduces destructive interferences between different paths in the graph 13. We note again that this is not possible in the transverse Ising model with random fields because the model is stoquastic and all the weights along the edges in the graph can be taken to be positive.
The XXZ chain does not have the stoquastic property so that, in principle, destructive interferences are possible. If we assume this to be the case, then a minimal model is the one discussed in Sec. A where the Stirling numbers of the second kind are replaced by modified Stirling numbers, see Eq. (A.4). I.e., our Lieb-Robinson bound will then be given by Eq. (B.1) with $S(k,j)\to\tilde{S}(k,j)$. We can then obtain a bound for information spreading by using
$\tilde{S}(k,j)\leq(-\text{e}^{1/\xi}\ln(1-\text{e}^{-1/\xi}))^{k}\equiv\alpha^{k}\,.$ (B.7)
Now the two sums can be performed and we find
$||[A_{0}(\tau),B_{\ell}(0)]||\leq\frac{4||A||\,||B||}{(\ell-1)!}\text{e}^{4CJ\alpha\tau}$ (B.8)
×$\times\left(\gamma(\ell,4CJ\alpha\tau)-2^{\ell-1}\text{e}^{-2CJ\alpha\tau}\gamma(\ell,2CJ\alpha\tau)\right)$
where $\gamma(\ell,x)$ is the lower incomplete Gamma function. For large $\ell$ we can approximate the Gamma function and find
$||[A_{0}(\tau),B_{\ell}(0)]||\lesssim\frac{2||A||\,||B||}{\ell!}(4CJ\alpha\tau)^{\ell}$ (B.9)
Using the Stirling approximation this means that the commutator is of order $\mathcal{O}(1)$ if $\tau\sim\ell/(4CJ\alpha\text{e})$. This means that the spreading is ballistic and since $\alpha=\alpha(\xi)$ increases with increasing localization length $\xi$, the velocity decreases with increasing disorder as expected. We stress once more though that we have not found any indications for such destructive interference—which would result in the modified Stirling numbers $\tilde{S}(k,j)$—to be present in the Heisenberg chain.
To summarize, an almost factorial growth of the commutator $[H,A_{0}]^{(k)}$, as expected in a generic quantum chaotic system, leads to exponential spreading while an exponential growth leads to ballistic spreading of information in Euclidean time.
Appendix C Off-diagonal matrix elements
We can also use our bounds on the growth of the commutator $[H,A]^{(k)}$ for a local operator $A$ to derive results for the asymptotic scaling of off-diagonal matrix elements of $A$ in the energy eigenbasis for a finite system.
We have
$\langle E_{i}|A(\tau)|E_{j}\rangle=A_{ij}\text{e}^{\omega\tau}$ (C.1)
[…]
$|A_{ij}|\leq||A(\tau)||\text{e}^{-\omega\tau}$ (C.2)
which is valid for any time $\tau$. We can thus minimize the r.h.s. with respect to $\tau$ to obtain an optimal bound.
C.1 General case
The general, quantum chaotic case was already analyzed in Ref. [30]. Here we provide a short summary. From Eq. (B.1) we see that
$||A(\tau)||\leq||A||\text{e}^{2q},\quad q=\text{e}^{2CJ\tau}-1\,.$ (C.3)
Minimizing the function $e^{2q}\text{e}^{-\omega\tau}$ w.r.t. $\tau$ one finds
$|A_{ij}|\leq||A||\left\{\begin{array}[]{cc}\exp(2\tilde{\omega}(1-\ln\tilde{\omega})-2)&\quad\tilde{\omega}>1\\ 1&\quad\tilde{\omega}\leq 1\end{array}\right.$ (C.4)
with $\tilde{\omega}=\omega/(4CJ)$. This means that matrix elements are in general non-zero even for extensive energy differences $\omega$. Asymptotically, these matrix elements decay faster than exponential.
C.2 Non-interacting case
In the non-interacting case, energy differences $\omega$ cannot be extensive. A more precise bound can be obtained by using again Eq. (C.2) together with the bound
$||A(\tau)||\leq\sum_{k}\frac{\tau^{k}}{k!}\underbrace{||[H,A]^{(k)}||}_{\leq||A||(8CJ)^{k}}$ (C.5)
where we have used Eq. (II.17). This leads to the upper bound
$|A_{ij}|\leq||A||\exp[(8cJ-\omega)\tau]\,.$ (C.6)
Minimizing the r.h.s. we thus find $|A_{ij}|=0$ if $\omega>8CJ$ and the generic bound $|A_{ij}|\leq||A||$ if $\omega\leq 8CJ$.
In the disordered, non-interacting case a sharp cut-off will remain. To be concrete, we can for example consider the $\sigma^{z}$ operator in the Anderson case. Using Eq. (C.5) together with the bound (III.15) we find
$|\sigma^{z}_{ij}|=0,$ $\quad\omega>8(D+1)$
$|\sigma^{z}_{ij}|\leq\frac{1}{D+1},$ $\quad\omega\leq 8(D+1)$
In addition to the sharp cut-off in energy, we also find a non-trivial bound for the magnitude of the matrix elements below that cutoff in this case which is consistent with the fact that the eigenstates of the Anderson model are eigenstates of $\sigma^{z}$ for $D\to\infty$. Off-diagonal matrix elements therefore vanish entirely in this limit.
C.3 Many-body localization
As for the Lieb-Robinson bounds, the general result (C.4) will still apply qualitatively as long as $[H,A]^{(k)}$ does grow almost factorially. We know this to be the case for the transverse Ising model with random fields and found strong evidence that this is also true in the disordered Heisenberg model.
We can, however, check again what would happen if there are destructive interferences between different paths in the commutator graph, leading to our model with the modified Stirling numbers $\tilde{S}(k,j)$, see Eq. (A.6). Using the bound (C.5) with $||[H,A]^{(k)}||\leq(2CJ)^{k}||A||\sum^{\prime}_{I_{1},\dots,I_{k}}\leq 2(4CJ\alpha)^{k}||A||$ with $\alpha=-\text{e}^{1/\xi}\ln(1-\text{e}^{-1/\xi})$ we find
$||A(\tau)||\leq 2||A||\exp(4CJ\alpha\tau)$ (C.8)
This then implies
$|A_{ij}|=0,$ $\quad\omega>4CJ\alpha$
$|A_{ij}|\leq||A||,$ $\quad\omega\leq 4CJ\alpha\,.$
It is instructive to consider the limits of weak and strong disorder. For weak disorder, the norm of the local Hamiltonian scales as $||h||\sim CJ\sim 1$ if we set the exchange couplings to be of order $\mathcal{O}(1)$. For the parameter $\alpha$ we find that $\alpha\sim\ln\xi$ for $\xi\gg 1$. This means that there is no cut-off for $\xi\to\infty$; extensive energy differences are allowed, qualitatively consistent with the general bound (C.4). For very strong disorder, on the other hand, $||h||\sim CJ\sim D$ where $D$ is the disorder strength. The parameter $\alpha$ for $\xi\ll 1$ becomes $\alpha\sim 1$. For very strong disorder in a finite system the prediction of this model is therefore that a sharp cut-off of order $4D$ exists. This is consistent with the infinite disorder limit where the eigenstates are Ising states and local operators can at most couple states which differ by a local spin flip.
To summarize, an almost factorial growth of the commutator $[H,A]^{(k)}$ implies that matrix elements are non-zero even for extensive energy differences but that they decay faster than exponential with the magnitude of the energy difference. For an exponential growth, on the other hand, there is a sharp cut-off and matrix elements for states with an energy difference larger than this cut-off are zero.
Appendix D Graphs
@@ D.1 No disorder
$p(m,k)$ $=$ $\sum_{i_{1}+\ldots+i_{m}=\frac{k}{2}}\;\prod_{j=1}^{m}C(i_{j}-1)$ (D.2)
@@
$s^{(1)}_{1}(k)=4^{k}\sum_{m=1}^{\frac{k}{2}}2^{m}\frac{m}{k-m}{{k-m}\choose{\frac{k}{2}}}=4^{k}{{k}\choose{\frac{k}{2}}}.$ (D.3)
@@
$s^{(1)}_{2}(k)=\frac{1}{4}s_{1}(k+1)=2^{2k+1}{{k}\choose{\frac{k-1}{2}}}\,.$ (D.4)
@@
$s^{(1)}_{l}(k)=\frac{1}{4}s_{l-1}(k+1)-s_{l-2}(k).$ (D.6)
Simple inductive arguments show that the solution to Eq. (D.6) with base cases  (D.4) and  (D.1) is indeed our previous result  (III.1.1).
D.2 Disordered case
@@
$||[H,\sigma^{z}_{0}]||_{1}\leq 8^{k}(D+1)^{k-1}.$ (D.7)
This is the result Eq. (III.15). The bound is exponential as expected, with an increasing disorder strength corresponding to an increased growth rate. We also see that setting $D=0$ gives us $8^{k}$, the exact result in the case without disorder.
@@
where ${}_{2}F_{1}$ is the hypergeometric function. Note that this result is derived for even $k$. For $k$ odd the support will be zero; one can infer this from Fig. 4 since there is no possible walk on the graph beginning and ending at the top node with an odd number of steps. The upper bound in Eq. (D.2) only arises from the fact that $D$ is an upper bound on the local magnetic fields; the enumerative calculations are exact. Similar numerical calculations show that the support on $l>1$ sites is bounded from above by a hypergeometric function as well but with a different leading term
@@
This coincides with Eq. (III.18). As mentioned in the main text, there is a maximum in this ratio which shifts to larger $l$ with increasing $k$, which is not expected for a localized model. This is likely explained by the looseness of the bound Eq. (D.2); while it is correct for the leading term in $D$, the actual prefactors for terms of decreasing order in $D$ become increasingly smaller than those in the given bound. We expect that a formula replacing Eq. (D.2) in which the enumerative calculations are exact and the upper bound only comes from $D$ would not possess this unexpected behaviour, although we have been unable to derive such a formula. We nevertheless see the expected result that with $k$ held constant we have an exponential decay of the ratio of support for increasing $l$, and that increasing $D$ corresponds to an increasing rate of decay.